OpenStudy (anonymous):

help please? how can −4x−2=−x+3 be graphed to find the point of intersection?

OpenStudy (anonymous):

@phi

Directrix (directrix):

I think this is supposed to be a system of equations: y = −4x−2 y = −x+3

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Is this how the problem appears in your text? >> −4x−2=−x+3

OpenStudy (anonymous):

yes, this is the whole problem: The navigator has to program the ship’s computer so that it will be prepared to intercept the satellite. To program it, the satelite’s path is set equal to the ship’s. The programmed equation is –4x – 2 = –x + 3. Explain to the navigator how the programmed equation can be graphed to find the point of intersection. Use complete sentences. @Directrix

Directrix (directrix):

My interpretation of the problem is to solve this system of equations by graphing to get the point of intersection of the ship and the satellite. No, I am not sure of this. y = −4x−2 y = −x+3